# Adaptive path planning for efficient object search by UAVs in agricultural fields

### v1

@@ 4.3 Impact of localization errors
Experiment 3 showed that the adaptive planner better deals with localization errors than the baseline coverage planner, till a position error of around 3 cm and an orientation error of around 1 degree. Because the localization error depends on the altitude, the low-altitude inspections decreased the localization error. Using a UAV equipped with RTK-GNNS should be accurate enough, indicating the practical usability of the planner. Additionally, the required accuracy of the localization also depends on the application. For instance, in a weed detection application, the resulting map may be used to sent a weeding robot to a specific spot in the field. In that case, the required localization accuracy may be lower, since it is only needed to guide the weeding robot until the weeds are within its field-of-view.
@@ 4.4 Validity of the offline simulation method
The adaptive path planner was evaluated using a simulated flight over the orthomosaics created based on real drone images. Since orthomosaics are corrected for perspective distortion, the resulting images may simplify real-world images slightly. However, since the plants were much smaller than the flight altitude of the drone, taking snapshots of parts of the orthomosaic resulted in a similar perspective as during a real-world flight. The F1-scores on the real images in experiment 1 at 12m altitude were comparable with the F1-scores from the coverage planner at 12m altitude from the simulated images in experiment 2, 3, and 4. Additionally, the typical object detection performance in agriculture ranges from 0.7 to 0.9 in F1-score (Rai et al., 2023; Ruigrok et al., 2023; Rehman et al., 2024), which is comparable with the detection performance observed in our experiments. This indicates that the orthomosaic provides an accurate simulation of images taken in a real-world flight. Future work will focus on applying the path planner in the real-world.
@@ 4.5 Operational opportunities for an adaptive path planner
By increasing the flight speed of the UAV at higher altitudes, the advantages of an adaptive planner could be larger. At a higher altitude, the field-of-view of the UAV is larger. Since the flight speed depends on the motion blur, it is possible to increase the flight speed at larger altitudes (Seifert et al., 2019). This will result in a higher efficiency of the adaptive planner compared to a low-altitude coverage planner than showed in this paper.
In this work, we used the same camera at high and low altitude, however implementing multi-modal sensing could further increase the efficiency of the method. For example, detection of biotic stress of plants caused by infections can be done using thermal images (Pineda et al., 2020), although these cameras have a low spatial resolution limiting their field-of-view. By combining a thermal camera with the RGB camera using the adaptive path planner, suspicious areas can be identified at high altitude using a high-resolution RGB camera, allowing for targeted closer inspection with a thermal camera at those specific spots.
@@ 5 Conclusion
In this study, we presented an adaptive path planning method that combines a high-altitude coverage path with low-altitude inspections based on the detection certainty. The path planner was evaluated on real-world orthomosaics derived from drone-based RGB images. Different detection certainty measures were compared, showing the standard YOLOv8 confidence value to be the best in differentiating true and false positive detections, thus most fit to be used in the adaptive path planner to determine on the low-altitude inspection of detected, but uncertain plants. When the plants were distributed in clusters, the adaptive planner yielded a 37% shorter flight path compared to a low-altitude coverage planner at the cost of only 2% lower F1-score. When the objects were uniformly distributed in the field, the adaptive path planner showed a 6% shorter flight path at the cost of a 2% lower F1-score. The adaptive path planner could better deal with errors in localization of the UAV than a coverage planner by performing low-altitude inspections. When the objects are uniformly distributed, a coverage path was more efficient, except when there were less than 80-120 objects per hectare in the field. However, when the objects were non-uniformly distributed, the adaptive path planner was faster than a low-altitude coverage path till 250-300 objects per hectare. In conclusion, the presented adaptive path planner allowed to find non-uniformly distributed objects in a field faster than a coverage path planner and resulted in a comparable detection accuracy.

### v2

@@ 4.4 Impact of localization errors
Experiment 3 showed that the adaptive planner better deals with localization errors than the baseline coverage planner, till a position error of around 3 cm and an orientation error of around 1 degree. Because the localization error depends on the altitude, the low-altitude inspections decreased the localization error. Using a UAV equipped with RTK-GNSS should be accurate enough, indicating the practical usability of the planner. Additionally, the required accuracy of the localization also depends on the application. For instance, in a weed detection application, the resulting map may be used to send a weeding robot to a specific spot in the field. In that case, the required localization accuracy may be lower, since it is only needed to guide the weeding robot until the weeds are within its field-of-view.
@@ 4.5 Validity of the offline simulation method
The adaptive path planner was evaluated using a simulated flight over the orthomosaics created based on real drone images. Since orthomosaics are corrected for perspective distortion, the resulting images may simplify real-world images slightly. However, since the plants were much smaller than the flight altitude of the drone, taking snapshots of parts of the orthomosaic resulted in a similar perspective as during a real-world flight. The F1-scores on the real images in experiment 1 at 12m altitude were comparable with the F1-scores from the coverage planner at 12m altitude from the simulated images in experiments 2, 3, and 4. Additionally, the typical object detection performance in agriculture ranges from 0.7 to 0.9 in F1-score (Rai et al., 2023; Ruigrok et al., 2023; Rehman et al., 2024), which is comparable with the detection performance observed in our experiments. This indicates that the orthomosaic provides an accurate simulation of images taken in a real-world flight. Future work will focus on applying the path planner in the real world.
@@ 4.6 Operational opportunities for an adaptive path planner
By increasing the flight speed of the UAV at higher altitudes, the advantages of an adaptive planner could be larger. At a higher altitude, the field-of-view of the UAV is larger. Since the flight speed depends on the motion blur, it is possible to increase the flight speed at larger altitudes (Seifert et al., 2019). This will result in a higher efficiency of the adaptive planner compared to a low-altitude coverage planner, as shown in this paper.
In this work, we used the same camera at high and low altitude; however, implementing multi-modal sensing could further increase the efficiency of the method. For example, detection of biotic stress of plants caused by infections can be done using thermal images (Pineda et al., 2020), although these cameras have a low spatial resolution, limiting their field-of-view. By combining a thermal camera with the RGB camera using the adaptive path planner, suspicious areas can be identified at high altitude using a high-resolution RGB camera, allowing for targeted closer inspection with a thermal camera at those specific spots.
@@ 5 Conclusion
In this study, we presented an adaptive path planning method that combines a high-altitude coverage path with low-altitude inspections based on the detection certainty. The path planner was evaluated on real-world orthomosaics derived from drone-based RGB images. Different detection certainty measures were compared, showing the standard YOLOv8 confidence value to be the best in differentiating true and false positive detections, thus most fit to be used in the adaptive path planner to determine the low-altitude inspection of detected, but uncertain plants. When the plants were distributed in clusters, the adaptive planner yielded a 37% shorter flight path compared to a low-altitude coverage planner at the cost of only 2% lower F1-score. When the objects were uniformly distributed in the field, the adaptive path planner showed a 6% shorter flight path at the cost of a 2% lower F1-score. The adaptive path planner could better deal with errors in the UAV’s localization than a coverage planner by performing low-altitude inspections. When the objects were uniformly distributed, a coverage path was more efficient, except when there were fewer than 80-120 objects per hectare in the field. However, when the objects were non-uniformly distributed, the adaptive path planner was faster than a low-altitude coverage path till 250-300 objects per hectare. In conclusion, the presented adaptive path planner allowed finding non-uniformly distributed objects in a field faster than a coverage path planner and resulted in a comparable detection accuracy.